OpenStudy (anonymous):

Which of the following sums would be under the radical symbol to find the distance between the points (7, -1) and (-8, -9)? 1^2+8^2?

OpenStudy (ash2326):

Distance formula between two points with coordinates (x1, y1) and (x2, y2) is given as \[\sqrt{(x2-x1)^2+(y2-y1)^2}\] here (x1, y1)= (7, -1) and (x2, y2)=(-8, -9)

OpenStudy (anonymous):

1^2+10^2??

OpenStudy (anonymous):

You are ignoring the negative sign..

OpenStudy (anonymous):

\(x_2 = -8\) \(x_ 1 = 7\) So: \(x_2 - x_1 = -8 - (7) = ??\)

OpenStudy (anonymous):

ooh 15^2+8^2

OpenStudy (ash2326):

yes:)

OpenStudy (anonymous):

Great.. Take care of negative signs..